HDR10 not working since today's update

Installed the big April update yesterday, had another update notice today. HDR is not working since then, just a “no signal” notice on my lg oled b6. The new “Force 422 subsampling” makes it work again but all the colors get wonky.

Really? I have the exact same TV model and the Vero 4K does HDR just as fine as before. What’s the content of your /sys/class/amhdmitx/amhdmitx0/attr? See this thread for example.

It says 444,10bit. Did you get another update notice today aswell? Everything worked fine yesterday… here is my log, no idea how to work with it: http://paste.osmc.tv/eyesediqok

If you are setting attr ‘manually’ eg by a line in rc.local that should no longer be necessary. Setting 422 may be necessary but shouldn’t make colours wonky. If you still have trouble, please post logs and define ‘wonky’.

1 Like

Ok I’ve found the solution. Since the newest update there is some sort of ‘allowed resolution to switch to’ setting, where I had previously selected all resolutions. Seems like the system got confused or something but selecting only the couple ones I need fixed my issues.

1 Like